OpenStudy (anonymous):

ok what if it has the > with a line under it like this one ..n + 14 >16 the line is under the >

OpenStudy (mertsj):

That means greater than or equal to.

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} \ge&\le\\ \uparrow &\uparrow \\ \textit{greater than or equal}& \textit{less than or equal} \end{array}\)

OpenStudy (anonymous):

would the answer be n < 2 but with the line under it

OpenStudy (jdoe0001):

\(\large \bf n+14\ge 16\implies n\cancel{+14-14}\ge 16-14\implies n\ge 2\)

OpenStudy (anonymous):

thank u that is the answer i got

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

what about b/14 < 16

OpenStudy (anonymous):

is the answer b<224

OpenStudy (jdoe0001):

yeap \(\bf \cfrac{b}{14} < 16\implies \cancel{14}\times \cfrac{b}{\cancel{14}} < 14\times 16\)